-- KrystsinaPetukhova - 2018-09-26

Timing issues observed during 2018 data-taking

channel problem affected run and LBs actions channel overall status supporting plots
EBA51 ch00-23 unstable timing for half a module
  • happens during the run
  • related to LVPS power-cycling of the EBA51 module (typically)
  • change in mean time happens simultaneously for each channel
  • change in mean time is approximately the same for all digitizers

two periods with unstable timing in 2018, up to now
  • few from 348251-349693 (frequent reconfigurations)
  • some from 355109-357193 (multiple automatic reconfigurations, manual power cycle)

  • periods with time jump will be corrected for 2018 data reprocessing
  • LBs with very unstable timing to be flagged with BAD_TIMING

  • The reason was found out to be unstable 3V output current on TTC - interface card to be replaced during the maintenance
  • discarding >70% events from 18-Oct-2018

EBA51 ch00 06 run348354 all digi.png
Timing instability in run 348354, all channels in digitizer (and generally ch00-23) simultaneously affected.
EBA51 ch00 ch06 run357193 all digi.png
Timing instability in EBA51 ch00-23 during the run 357193.
EBA51 ch00 time offset about (+3) ns
  • occurred during 360161, since LB=296, directly following power cycle of EBA51
  • up to now

time constant to be corrected for reprocessing
  • provided good pulse shape
  • no change in reconstructed energy

no response in LG in laser and CIS calibration runs since 350161
LBA51 ch00 shift.png
Timing shift in EBA51 ch00 occurred during the run 360161: at LB=298 (power cycle of the module) and LB=331 (LVPS trip) .
EBA12 ch 20 fraction of events has 2x bunch-space offset (-50) ns
  • a fraction of events with (-50) ns offset is typically below 1%
  • mean time in physics seem sot be not affected

almost each run since 348403 when number of colliding bunches increased from 8 to 62 The issue concern all DMU 6 (ch 18 and 19 are absent). Out-of-time events could be separated in laser (provided the signal is known). The cut which will help to distinguish offset events in physics data is under development. Decided not to mask the channel
  • provided relatively small fraction of affected events
  • no significant impact of physics data

OK on DMU level
EBA12 ch20 run340697.png
Fraction of events registered by EBA12 ch20 has (-50) ns (run 357193 shown).
PS EBA12 ch 20 run351223 lb0 lb1080.png
Pulse shape in EBA12 ch20 has (-50) ns (run 351223, all LBs, shown).
EBA40 ch 39-41 fraction of events has 2x bunch-space offset (+25) ns
  • a fraction of events with (+25) ns offset is typically below 1%
  • mean time in physics seem sot be not affected

almost each run since 348403 when number of colliding bunches increased from 8 to 62 The issue concern all DMU 13 (ch 18 and 19 are absent). Out-of-time events could be separated in laser (provided the signal is known). The cut which will help to distinguish offset events in physics data is under development. Decided not to mask the channel
  • provided relatively small fraction of affected events
  • no significant impact of physics data

OK on DMU level
EBA40 ch36 run340697.png
Fraction of events registered by EBA40 ch39-41 (DMU5) has (+25) ns (run 340697 shown).
EBA27 ch01 time offset about (-4) ns
  • occurred during 357679, since LB=550, directly following automatic power cycle of EBA27 as over-current
  • up to now

time constant to be corrected for reprocessing
  • provided slight undershoot in pulse shape in HG and not reasonable pulse shape in LG
  • change of reconstructed in HG energy by 20-25% (from 1.3 pC to 1.6 pC) although eFit distribution is Gaussian-like

no CIS response in both gains since 357679 (supposedly burned 3-in-1 card)
EBA27 ch01 run357713.png
Timing offset in EBA27 ch01 (run 357713 shown).
PulseShape EBA27 ch01 357713 lb0 -1.png
Pulse shape in EBA27 ch01 (run 357713 shown).
EBA53 ch01 sporadic offset (+3) ns seen in laser-in-gap events
  • happens during the run
  • concerns only one channel -- ch01

  • some from 349014-350751
  • 354309, 355261, 355848
  • some from 357409-359191
  • 360293
  • mean time in physics data is not seemingly affected

zeros in HG followed by switch to LG where signal is mis-reconstructed (bad pulse shape in LG)
  • increasing pedestal for this channel to be performed on software level
  • to be flagged with BAD_TIMING

LG ADC masked (unspecified bad). The channel is known for problems with gain switch.
EBA53 ch01 LMT.png
Timing instability in EBA53 ch01.
EBA53 ch01 PS.png
Pulse shape in EBA53 ch01, during LBs with affected timing.
EBA16 ch00 sporadic offset (-3) ns
  • happens during the run
  • concerns only one channel -- ch00

from 349534 up to now
  • observed from run to run
  • mean time in physics data is about (-1) ns

zeros in HG followed by switch to LG where signal is mis-reconstructed (bad pulse shape in LG)
  • increasing pedestal for this channel to be performed on software level
  • to be flagged with BAD_TIMING

since (335237,0) HG marked as no CIS calibration. The channel is known for problems with gain switch (supposedly burnt 3in1 card).
EBA16 ch00 run355468.png
Timing instability in EBA16 ch00-23.
EBA08 ch07 time offset about (-4) ns
  • occurred during 359191, since LB=922, directly following automatic power cycle of EBA08 as over-current
  • up to now

set bad timing from (359191, 922)
  • provided distorted pulse shape in HG and reasonable pulse shape in LG
  • change of reconstructed in HG energy by 20-25% (from 3.2 pC to 4 pC) and eFit with pultiple-peak structure

CIS and laser calibrations
  • no CIS response in both gains since 360833 (supposedly 3-in-1 card problem)
  • laser constant drift since 360833

EBA08 ch07 offset run359286.png
Timing in EBA08 ch07 during the run 359286.
PulseShape EBA07 ch07 run 360560 lb0 -1.png
Pulse shape in EBA08 ch07 (run 360560 shown).
EBA08 ch07 run 359279 circled.png
Time offset in 359279 in physics data.
EBC57 ch11 time offset of varying value
  • observed only on a single channel (ch11 in extended barrel - A13)
  • seen on laser-in-gap events and physics data
  • time offset happens out of run, first LB already has the offset

some runs from 349114-351455
  • recovered after power cycle during 351455
  • seems to be stable in 351550 and onwards

  • pulse shape is OK
  • to be corrected for reprocessing

calibration (observed throughout 2018 data-taking, not only in the runs with affected timing)
  • mono-gain CIS: 100% SB5-9 in LG
  • x2 larger RMS of amplitude in pedestal runs

EBC57 ch11 runN349335.png
Time offset in EBC57 ch11 in run 349335.
run 351062.png
Time offset in EBC57 ch11 in run 351062 (physics data).
pulse shape EBC57 ch11 run351062.png
Pulse shape in EBC57 ch11 in run 351062.
EBA19 ch00 mean time changes (+- 10 ns roughly)
  • seen in laser-in-gap events and physics data
  • following HV change and change of data-taking conditions
  • since 348836 (+2.0) ns in laser-in-gap-events
  • could vary in physics (due to non-gaussian shape of time distribution and/or low statistics runs). Please, report if the offset is above 10 ns in physics, and time in laser changes (about (+2.0) ns now)

two periods (up to now):
  • 348403-348836 (changes of HV applied)
  • 354124-355468 (vdM, beta and low-mu runs)

pulse shape in LG is not good. Overflow in HG -> switch to LG laser LG calibration run: very large reconstructed response (evolving: 900+-35 pC at the beginning of data-taking, 4500+-180 pC in September 2018)
EBA19 ch0 runN348836.png
Time offset in EBA19 ch00 in run 348836.
EBA19 ch00 pulse shape.png
Pulse shape in EBA19 ch00.
EBC25 ch15-17 time offsets in a few LB seen in laser-in-gap events
  • no effect on physics (several LBs with time offset)

throughout data-taking
  • run 348895, 350803, 351270, 351698, 352030, 352494 , 354494, 355224, 355563, 355599, 357293, 357539, 358175, 358300, 358325, 358333, 358577, 358615, 359471, 359472, 359717
  • may happen another time. Please, report about it.

pulse shape is not good, even in runs with well reconstructed timing
  • candidate for masking , DMU replacement

  • large noise in mono-gain CIS and laser calibration runs (in HG x30, in LG ~x7)
  • saturation 1% in laser HG run (could be indication of wrong switch circuit)

EBC25 ch15 17 run362345.png
Time in EBC25 ch15-17 (run 362345).
EBC25 ch15 pulse shape.png
Pulse shape in EBC25 ch15.
EBC38 ch01 time offset (-4) ns
  • occurred during the run

runs 352436-353907 (laser-in-gap)
  • since (354021, 0) EBC38 ch01 LG flagged with bad timing
  • since (354477, 0) EBC38 ch01 LG no CIS calibration

zeros in HG (in last sample) leading to switch to LG
  • pulse shape distorted
  • energy affected by 10-20%
  • flagged with bad timing

no CIS calibration
EBC38 ch01 run352436.png
Time in EBC38 ch01 (run 352436), after offset appeared.
PulseShape EBC38 ch01 run352436.png
Pulse shape in EBC38 ch01, after the offset appeared.
EBA10 ch08 time offset (-3.5) ns
  • occurred during the run
  • affected after overcurrent and automatic power cycle as over-current (LBs affected due to the trip: 442 - 470)

since (363710, 480)
  • distorted pulse shape
  • undershoot in pulse shape
  • reconstructed energy is overestimated w.r.t. its usual value: ~25%
  • decision: correct timing for reprocessing (?)

  • monoCIS LG: no response in LG, noise-like amplitude in HG
  • monoCIS HG: no amplitude reconstructed in HG
  • laser LG: 10% zeros, time offset
  • laser HG: response in both gains

EBA10 ch08 run363710.png
Time shift in EBA10 ch00 happened during the run 363710.
EBA10 ch08 run363830 pulse shape.png
Pulse shape in EBC10 ch08, in the run 363830, after the offset appeared.
EBC50 ch01 time offset (-3.5) ns
  • occurred during the run
  • affected after automatic reconfiguration (due to overcurrent) and trip

since (363979,500)
  • distorted pulse shape with undershoot
  • reconstructed energy is over-estimated by ~25%
  • decision: correct timing for reprocessing (?)

  • monoCIS LG: no response in LG, noise-like amplitude in HG
  • monoCIS HG: no amplitude reconstructed in HG
  • laser: time offset, indication of zeros (~16%)

EBC50 ch08 run363979.png
Time shift in EBA50 ch08 happened during the run 363979.
EBC50 ch08 run364030 pulse shape.png
Pulse shape in EBC50 ch08, after the offset appeared, in the run 364030.
EBC40 ch16 time offset (-3.5) ns
  • occurred on the fly
  • neither LVPS trip nor power cycle at that moment
  • similar pattern is seen for channels in odd-numbered DMUs: 1, 3, 5, 7; time offset is present for all channels, but to smaller extent

since (364214, 492)
  • undershoot in samples [0..2] in some events
  • double peak in reconstructed energy
  • decision -- to mask this channel
  • the channel is already masked in runs 364214, 364292
  • will be observed in heavy ions collisions

  • ch21,22 are dead in monoCIS (LG and HG) and laser

EBC40 ch16 run364214.png
Time shift in EBC40 ch16 occurred during the run 364214.
EBC40 ch16 run364292 pulse shape after.png
Pulse shape in EBC40 ch16, in the run 364292, following one after the offset appeared.
EBC40 ch16 run364214 time energy corr bad.png
Reconstructed energy-time scattering plot: events with ~(-4) ns have 25% higher energy reco'ed.
$\mathcal{S}_{E_{\mathrm{T}}^{\mathrm{miss}}}$
Topic attachments
I Attachment History Action Size Date Who Comment
PNGpng EBA08_ch07_offset_run359286.png r1 manage 53.2 K 2018-09-26 - 22:15 KrystsinaPetukhova  
PNGpng EBA08_ch07_run_359279_circled.png r1 manage 40.5 K 2018-09-26 - 22:15 KrystsinaPetukhova  
PNGpng EBA10_ch08_run363710.png r1 manage 16.6 K 2018-10-29 - 14:25 KrystsinaPetukhova  
PDFpdf EBA10_ch08_run363830_pulse_shape.pdf r1 manage 43.8 K 2018-11-01 - 15:20 KrystsinaPetukhova  
PNGpng EBA10_ch08_run363830_pulse_shape.png r1 manage 164.1 K 2018-11-01 - 15:42 KrystsinaPetukhova  
PNGpng EBA12_ch20_run340697.png r1 manage 12.3 K 2018-09-26 - 22:15 KrystsinaPetukhova  
PNGpng EBA16_ch00_run355468.png r1 manage 61.8 K 2018-09-26 - 22:15 KrystsinaPetukhova  
PNGpng EBA19_ch00_pulse_shape.png r1 manage 38.0 K 2018-10-04 - 13:21 KrystsinaPetukhova  
PNGpng EBA19_ch0_runN348836.png r1 manage 56.1 K 2018-10-04 - 13:21 KrystsinaPetukhova  
PNGpng EBA27_ch01_run357713.png r1 manage 60.7 K 2018-09-26 - 22:15 KrystsinaPetukhova  
PNGpng EBA40_ch36_run340697.png r1 manage 33.8 K 2018-09-26 - 22:15 KrystsinaPetukhova  
PNGpng EBA51_ch00_06_run348354_all_digi.png r1 manage 9.6 K 2018-09-26 - 22:15 KrystsinaPetukhova  
PNGpng EBA51_ch00_ch06_run357193_all_digi.png r1 manage 10.7 K 2018-09-26 - 22:15 KrystsinaPetukhova  
PNGpng EBA53_ch01_LMT.png r1 manage 8.6 K 2018-09-26 - 22:15 KrystsinaPetukhova  
PNGpng EBA53_ch01_PS.png r1 manage 42.7 K 2018-09-26 - 22:15 KrystsinaPetukhova  
PNGpng EBA53_ch0_run349114.png r1 manage 34.4 K 2018-09-26 - 22:15 KrystsinaPetukhova  
PNGpng EBC25_ch15_17_run362345.png r1 manage 31.6 K 2018-10-09 - 14:44 KrystsinaPetukhova  
PNGpng EBC25_ch15_pulse_shape.png r1 manage 34.6 K 2018-10-04 - 13:40 KrystsinaPetukhova  
PNGpng EBC38_ch01_run352436.png r1 manage 8.0 K 2018-10-04 - 15:16 KrystsinaPetukhova  
PNGpng EBC40_ch16_run364214.png r1 manage 16.6 K 2018-10-29 - 14:25 KrystsinaPetukhova  
PNGpng EBC40_ch16_run364214_time_energy_corr_bad.png r1 manage 9.5 K 2018-10-29 - 14:25 KrystsinaPetukhova  
PNGpng EBC40_ch16_run364292_pulse_shape_after.png r1 manage 34.9 K 2018-10-29 - 14:25 KrystsinaPetukhova  
PNGpng EBC50_ch08_run363979.png r1 manage 18.0 K 2018-10-29 - 14:25 KrystsinaPetukhova  
PNGpng EBC50_ch08_run364030_pulse_shape.png r1 manage 169.3 K 2018-11-01 - 15:22 KrystsinaPetukhova  
PNGpng EBC57_ch11_runN349335.png r1 manage 44.5 K 2018-10-04 - 13:21 KrystsinaPetukhova  
PNGpng LBA51_ch00_shift.png r1 manage 16.4 K 2018-09-26 - 22:15 KrystsinaPetukhova  
PNGpng PS_EBA12_ch_20_run351223_lb0_lb1080.png r1 manage 36.0 K 2018-09-26 - 22:15 KrystsinaPetukhova  
PNGpng PulseShape_EBA07_ch07_run_360560_lb0_-1.png r1 manage 38.0 K 2018-09-26 - 22:15 KrystsinaPetukhova  
PNGpng PulseShape_EBA27_ch01_357713_lb0_-1.png r1 manage 35.8 K 2018-09-26 - 22:15 KrystsinaPetukhova  
PNGpng PulseShape_EBC38_ch01_run352436.png r1 manage 57.1 K 2018-10-04 - 15:16 KrystsinaPetukhova  
PNGpng pulse_shape_EBC57_ch11_run351062.png r1 manage 32.3 K 2018-10-04 - 13:21 KrystsinaPetukhova  
PNGpng run_351062.png r1 manage 29.8 K 2018-10-04 - 13:21 KrystsinaPetukhova  
Edit | Attach | Watch | Print version | History: r21 < r20 < r19 < r18 < r17 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r21 - 2019-09-28 - KrystsinaPetukhova
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Sandbox All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback